# A Survey on Transformer Compression

### 05 Feb 2024 v1

@@ 3.1.3 Quantization for Transformer-Based Vision Models
Quantization-aware training. When compressing vision transformers to extremely low-bit precision, PTQ can not optimize the large quantization error with limited calibration images and suffers from significant performance reduction, thus QAT are urgently required for more accurate low-bit vision transformers. Q-ViT [48] finds that MSA and GELU are much sensitive to quantization, thus proposes a fully differentiable quantization method, which adopts head-wise bit-width and switch scale during quantization searching process. Quantformer [46] takes the self-attention rank into consideration, and proposes to maintain the consistency between quantized and full-precision vision transformers. Besides, Quantformer presents the group-wise strategy to quantize feature of patches in different dimensions, where each group adopts different quantization parameters and the extra computation overhead is negligible. Based on the observation that the sever performance degradation suffers from the quantized attention map, AFQ-ViT [47] designs an information rectification module and a distribution guided distillation during quantization training. The former one helps recover the distribution of attention maps with information entropy maximization in the inference, and the later one reduces the distribution variation with attention similarity loss in the backward. Similar with FQ-ViT [22], I-ViT [50] also explores the integer-only quantization scheme for ViTs. I-ViT designs Shiftmax, ShiftGELU and I-LayerNorm to replace the vanilla modules with bit-wise shift and integer matrix operations in the inference, achieving 3.724.11×3.72-4.11$\times$3.72 - 4.11 × acceleration compared to the floating-point model. OFQ [23] finds that weight oscillation causes unstable quantization-aware training and leads to sub-optimal results, and the oscillation comes from the learnable scale factor and quantized query and key in self-attention. To address that, OFQ proposes statistical weight quantization to improve quantization robustness, and freezes the weights with high confidence and calms the oscillating weights with confidence-guided annealing. For the query and key in the self-attention, OFQ presents query-key reparameterization to decouple the negative mutual-influence between quantized query and key weights oscillation. Similarly, VVTQ [56] analyzes ViT quantization from the perspective of variation, which indicates that the data variance within mini-batch are much harmful to quantization and slow down the training convergence. To reduce the impact of variations in the quantization-aware training, VVTQ proposes a multi-crop knowledge distillation-based quantization methodology, and introduces module-dependent quantization and oscillation-aware regularation to enhance the optimization process.
@@
We summarize the results of these PTQ and QAT methods in Table II, most schemes conduct PTQ on 8bit and 6-bit, there are server degradation when quantized to 4-bit. With the complete training data, QAT could push the ViT quantization to 4-bit and less, some methods even shows better results than floating point models, which proves that quantization-aware training could tap into the true accuracy of quantized models efficiently. And Figure 3 shows the latency results of ViT models using FasterTransformer [^1] https://github.com/NVIDIA/FasterTransformer.
@@ 3.1.4 Discussion
With the presence of extreme distributions and outliers, quantization for transformers is much more difficult than that for convolutional neural networks. To recover the performance of quantized models, various methods are proposed to address the quantization-unfriendly components in transformers. Specifically, For vision tasks, the existing methods optimizes the quantized transformers in three ways: remaining the self-attention rank, rectifying the distribution of extreme distribution and addressing the weight oscillation and data variation in quantization-aware training, as shown in Figure 2. And for natural language processing task, most schemes aims to process the outlier of weights and activations. Besides, the training overhead is usually unacceptable for large language models with billion-level parameters, thus quantizing module by module and conducting parameter-efficient fine-tune are much popular. However, when compressing to extremely low bit-width, the quantized transformers suffer significant performance degradation, which are far away from the floating-point models, thus how to build a more accurate and low-bit transformers is still a hard problem to be solved.
@@ 3.2.1 Overview of Knowledge Distillation
Knowledge distillation (KD) aims to train student networks by compressing [83, 84, 85] or transferring [87, 88, 86] knowledge from teacher networks. In this paper, our main focus lies on distillation methods proposed to achieve a compact student model while preserving satisfactory performance compared to a cumbersome teacher model. The student models typically have narrower and shallower architectures, making them more suitable for deployment on resource-limited systems.
We will first discuss two different kinds of knowledge distillation in the following: logits-based methods [87] which convey knowledge on the logits level and hint-based methods [89] which convey knowledge through intermediate features. To illustrate the objective of logits-based KD in classification tasks, we denote the logits output of the teacher and student network as $z^{\bf s},z^{\bf t}\in\mathbb{R}^{C}$, where $C$ represents the number of classes. Neural networks often generate class probabilities by applying a softmax function to the logits, converting them into probabilities $p^{\bf s},p^{\bf t}$ as follows:
@@ 3.2.2 KD for Transformer-Based Large Language Models
Logits-based KD. Tang et al. [90] employed knowledge distillation to compress BERT[91], a large language model, into a much lighter bidirectional long short-term memory network (BiLSTM) [92] for natural language processing (NLP) tasks. The distillation process aimed to minimize the mean squared error (MSE) loss between the student’s logits and the teacher’s logits. Following distillation, the shallow BiLSTM-based model achieved comparable results to ELMo[93], but with approximately 100$\times$ fewer parameters and a 15$\times$ faster inference speed. Similarly, DistillBERT [24] initialized the shallower student with teacher’s parameters, and minimized the soft target probabilities between the teacher and the student, also known as word-level KD. SeqKD [94] fine-tuned the student model on the sequence-level teacher-generated data. MixKD [95] extended the idea of encouraging the student to mimic the teacher’s logits to the linear interpolation of example pairs [96]. Theoretical analysis has shown that, under reasonable conditions, MixKD can effectively reduce the gap between generalization error and empirical error. Turc et al. [97] demonstrated the ongoing significance of pre-training, even when employing smaller architectures. They introduced a methodology called Pre-trained Distillation (PD), which begins with pre-trained compact student models. Subsequently, it explores the transfer of task knowledge from large fine-tuned models using conventional logits-based KD. In contrast, MINILLM [98] highlighted a limitation of conventional logits-based KD that minimize forward Kullback-Leibler divergence (KLD) in free-run generation. They found that such approaches can lead the student model to overestimate the low-probability regions of the teacher’s distribution. To address this issue, MINILLM proposed replacing the forward KLD objective (as shown in Eq. 7) with reverse KLD, i.e., $reverse~{}\mathrm{KLD}:=KL(p^{\bf s}||p^{\bf t})$, which is better suited for knowledge distillation on generative language models. GKD [99] identified a challenge in distilling knowledge into auto-regressive student models, which is the train-test distribution mismatch. Specifically, partial sequences encountered by the student during generation phase can be significantly different from the ones observed during the training phase. To address this issue, GKD trained the student on its self-generated output sequences by leveraging feedback (logits) from the teacher on such sequences. It also facilitated the integration of distillation with RL fine-tuning of language models. Huang et al. [100] further combined the hint-based KD with multitask in-context learning.
@@
Hint-based KD. Sun et al. [101] proposed patient knowledge distillation to help shallow BERT student learn from multiple intermediate features [102, 103, 104, 25] of the deep BERT teacher. Li et al. [105] utilized hints extracted from both intermediate hidden states and attention distributions to enhance the training of Non-AutoRegressive Translation (NART) models. The distilled NART models attained performance similar to a powerful LSTM-based AutoRegressive Translation (ART) baseline in various machine translation tasks. Notably, the distilled NART models exhibited a speed improvement of 17$\times$ over their ART counterparts. Similarly, Mukherjee et al. [106] employed MBERT [107] as a teacher model to guide the training of smaller BiLSTM models, resulting in a significant 51$\times$ speed improvement for batch inference. MobileBERT [103] implemented two objective functions to distill knowledge from a BERT teacher incorporated with inverted bottleneck, including attention distributions and hidden states, to a slimmed-down version of BERT as the student model. In addition to these objectives, TinyBERT [104] extended the distillation process by also transferring knowledge from the teacher’s logits to the student model. Wang et al. [25] proposed MINILM to train the student networks to mimic both the attention distributions and the scaled dot-product between values of teacher models. It retains more than 99% accuracy on GLUE benchmark tasks using 50% of the Transformer parameters and computations of the teacher model.
@@
Several recent studies [110, 111, 112, 113, 114, 7, 115, 8, 116] have showed the promising results in fine-tuning small models on the outputs generated by LLMs. Specifically, they first prompted a very large teacher model, such as GPT-3 with 175B parameters, to solve complex questions via chain-of-thought (CoT) reasoning [117]. Then, they used the generated instructions [7, 115], explanations [8] or rationales [110] to fine-tune a much smaller student. Furthermore, Zhu et al. [118] proposed PaD, which utilizes program-aided reasoning, such as an additional Python interpreter, to help student models overcome faulty steps in the CoT reasoning with automated error checking. Shridhar et al. [119] introduced Socratic CoT that trains a combination of two small distilled models: a problem decomposer and a subproblem solver. These models work together to decompose and solve complex problems effectively. Fu et al. [120] demonstrated that specializing the small student model’s abilities for the specific reasoning task, by transferring the knowledge in larger teacher model’s generic directions, yields promising results. Jiang et al. [7] identified a shortcoming in the aforementioned approaches, as they lacked the ability to incorporate feedback for identifying challenging instructions where the student model’s performance falls short. To address this limitation, they proposed a three-stage adversarial loop of distillation that incorporates feedback and addresses these challenging instructions.
Other KDs. Xu et al. [121] compressed the BERT model by progressive module replacing. The original model was first divided into multiple modules, and more compact substitutes are created for these modules. Then, they randomly substituted the original modules with their corresponding substitutes and train the compact modules without any additional loss functions.
Jha et al. [122] addressed the challenge of fitting both the student and teacher models, which consist of hundreds of parameters, into GPU memory, making traditional knowledge distillation infeasible. To overcome this limitation, they proposed a teacher-free task-agnostic distillation method which uses a truncated version (student) of the large model (teacher) for initialization [123] and then continuing the pre-training of student, without employing any distillation loss.
@@ 3.2.3 KD for Transformer-Based Vision Models
Logits-based KD. While conventional KD in vision tasks typically involves using soft labels [87, 125, 84] generated by a powerful teacher network, Touvron et al. [126] made a noteworthy discovery. Their experiments on the ImageNet benchmark revealed that employing hard labels, which are essentially the maximum of the score values predicted by the teacher model, led to superior results. Additionally, they introduced a novel distillation token into vision transformers (ViTs)[127]. TinyViT et al.[27] observed that smaller ViTs could benefit from larger teacher using massive pre-training data, i.e., pre-training (distilling) student on ImageNet-21k while fine-tuning student on ImageNet-1k. And they proposed TinyViT to save computational memory by storing data augmentation information and logits in advance for large teacher models. Ren et al. [128] argued that student accuracy is primarily influenced by the inductive bias of the teacher models rather than their accuracy. To address this, they introduced a cross-inductive bias distillation approach, which involves distilling the student model with multiple teachers exhibiting distinct architectural inductive biases. To enhance the utilization of the inductive bias in CNNs and offer more image-friendly guidance, CSKD [129] removed the pooling operation after the final feature map and treated the features at each position as individual samples. Subsequently, these position-based features were input into the classifier, leading to the generation of dense predictions for each corresponding position. The hard labels derived from CNN’s dense predictions were used as target labels, and cross-entropy served as the loss function for spatial knowledge transfer.
Hint-based KD. Chen et al. [130] introduced a two-stage DearKD framework. In the initial stage, knowledge distillation is employed from the intermediate layers of the CNN to help transformer-based student capture the inductive biases. Subsequently, the second stage trains the student without distillation. Hao et al. [28] proposed a fine-grained manifold distillation method to fully utilize the patch-level and batch-level information inside transformer-based architectures. Specifically, they manually selected intermediate layers from teacher-student pair and distilled the student via decoupled relation maps. ViTKD et al. [131] delved into the characteristics of feature maps in ViTs and formulated three practical guidelines for hint-based distillation in ViTs, namely: (i) generating hints is more effective than mimicking in deeper layers, (ii) distillation on the shallower layers is also suitable for ViTs with mimicking, and (iii) features from the feed-forward network (FFN) are better than those from the multi-head attention (MHA) for the distillation process.
@@ 3.2.4 Discussion
One of the strengths of ViT models is their ability to scale to high parametric complexity, but this requires substantial computational resources and comes at a high cost. KD can be used to transfer knowledge into more compact student models, but there are still challenges in the context of vision that require further research. The first challenge is related to training costs. Both logits-based and hint-based KD methods require large GPU memory for the distillation process. Vision tasks differ from NLP as the input for vision is in the form of images, which are significantly larger in size compared to the limited sequence length used in NLP tasks. During distillation, even though the teacher network doesn’t require backward propagation, each forward pass consumes a significant amount of GPU memory due to the activations of intermediate features. Therefore, finding ways to reduce training costs and shorten training times is an area that needs exploration. Additionally, choosing the most suitable teacher model for transformer-based students is an open question. For instance, in the context of ViTs, determining whether a CNN or a transformer teacher is more effective for the student model is a potential research avenue. Moreover, as proposed by [84], there is a necessity for designing and evaluating KD approaches within practical scenarios, moving away from the limitations of small-scale datasets. This ensures that KD methods are applicable and effective in real-world, large-scale settings.

### 07 Apr 2024 v2

@@ 3.1.3 Quantization for Transformer-Based Vision Models
Quantization-aware training. When compressing vision Transformers to extremely low-bit precision, PTQ can not optimize the large quantization error with limited calibration images and suffers from significant performance reduction. As such, QAT is urgently required for more accurate low-bit vision Transformers. Q-ViT [48] finds that MSA and GELU are highly sensitive to quantization, and so proposes a fully differentiable quantization method that adopts head-wise bit-width and switch scale during the quantization searching process. Quantformer [46] takes the self-attention rank into consideration, and proposes to maintain the consistency between quantized and full-precision vision Transformers. In addition, Quantformer presents the group-wise strategy to quantize feature of patches in different dimensions, where each group adopts different quantization parameters and the extra computation cost is negligible. Based on the observation that sever performance degradation suffers from the quantized attention map, AFQ-ViT [47] designs an information rectification module and a distribution guided distillation during quantization training. The former helps recover the distribution of attention maps with information entropy maximization in the inference, and the latter reduces the distribution variation with attention similarity loss in the backward. Similar with FQ-ViT [22], I-ViT [50] also explores the integer-only quantization scheme for ViTs. I-ViT designs Shiftmax, ShiftGELU and I-LayerNorm to replace the vanilla modules with bit-wise shift and integer matrix operations in the inference, achieving 3.724.11×3.72-4.11$\times$3.72 - 4.11 × acceleration compared to the floating-point model. OFQ [23] finds that weight oscillation causes unstable quantization-aware training and leads to sub-optimal results, and the oscillation comes from the learnable scale factor and quantized query and key in self-attention. To address that, OFQ proposes statistical weight quantization to improve quantization robustness, freezing the weights with high confidence and calming the oscillating weights with confidence-guided annealing. For the query and key in the self-attention, OFQ presents query-key reparameterization to decouple the negative mutual-influence between quantized query and key weights oscillation. Similarly, VVTQ [56] analyzes ViT quantization from the perspective of variation, which indicates that the data variance within mini-batch is much harmful to quantization and slow down the training convergence. To reduce the impact of variations in the quantization-aware training, VVTQ proposes a multi-crop knowledge distillation-based quantization methodology, and introduces module-dependent quantization and oscillation-aware regularization to enhance the optimization process.
@@
We summarize the results of these PTQ and QAT methods in Table II. Most schemes conduct PTQ on 8-bit and 6-bit, and there is server degradation when quantized to 4-bit. With the complete training data, QAT could push the ViT quantization to 4-bit and less. Some methods even show better results than floating point models, which prove that quantization-aware training could tap into the true accuracy of quantized models efficiently. Figure 3 shows the latency results of ViT models using FasterTransformer [^1] https://github.com/NVIDIA/FasterTransformer.
@@ 3.1.4 Discussion
With the presence of extreme distributions and outliers, quantization for Transformers is much more difficult than that for convolutional neural networks. To recover the performance of quantized models, various methods are proposed to address the quantization-unfriendly components in Transformers. Specifically, for vision tasks, the existing methods optimizes the quantized Transformers in three ways: retaining the self-attention rank, rectifying the distribution of extreme distribution, and addressing the weight oscillation and data variation in quantization-aware training, as shown in Figure 2. And for natural language processing task, most schemes aims to process the outlier of weights and activations. What’s more, the training overhead is usually unacceptable for large language models with billions of parameters, and therefore quantizing module by module and conducting parameter-efficient fine-tune are much popular. However, when compressing to extremely low bit-width, the quantized Transformers suffer significant performance degradation, and perform far worse than the floating-point models. As such, how to build a more accurate and low-bit Transformers is still a difficult problem to be solved.
@@ 3.2.1 Overview of Knowledge Distillation
Knowledge distillation (KD) aims to train student networks by compressing [83, 84, 85] or transferring [86] knowledge from teacher networks. In this paper, our main focus lies on distillation methods proposed to achieve a compact student model while preserving satisfactory performance compared to a cumbersome teacher model. The student models typically have narrower and shallower architectures, making them more suitable for deployment on resource-limited systems.
We will first discuss two different kinds of knowledge distillation in the following: logits-based methods [86, 87, 88, 89] which convey knowledge on the logits level and hint-based methods [90] which convey knowledge through intermediate features. To illustrate the objective of logits-based KD in classification tasks, we denote the logits output of the teacher and student network as $z^{\bf s},z^{\bf t}\in\mathbb{R}^{C}$, where $C$ represents the number of classes. Neural networks often generate class probabilities by applying a softmax function to the logits, converting them into probabilities $p^{\bf s},p^{\bf t}$ as follows:
@@ 3.2.2 KD for Transformer-Based Large Language Models
Logits-based KD. Tang et al. [88] employs knowledge distillation to compress BERT[91], a large language model, into a much lighter bidirectional long short-term memory network (BiLSTM) [92] for natural language processing (NLP) tasks. The distillation process aims to minimize the mean squared error (MSE) loss between the student’s logits and the teacher’s logits. Following distillation, the shallow BiLSTM-based model achieves comparable results to ELMo[93], but with approximately 99% fewer parameters and a 15$\times$ faster inference speed. Similarly, DistillBERT [24] initializes the shallower student with teacher’s parameters, and minimized the soft target probabilities between the teacher and the student, also known as word-level KD. SeqKD [94] fine-tunes the student model on the sequence-level teacher-generated data. MixKD [95] extends the idea of encouraging the student to mimic the teacher’s logits to the linear interpolation of example pairs [96]. Theoretical analysis has shown that, under reasonable conditions, MixKD can effectively reduce the gap between generalization error and empirical error. Turc et al. [97] demonstrates the ongoing significance of pre-training, even when employing smaller architectures. They introduce a methodology called Pre-trained Distillation (PD), which begins with pre-trained compact student models. Subsequently, it explores the transfer of task knowledge from large fine-tuned models using conventional logits-based KD. In contrast, MINILLM [98] highlights a limitation of conventional logits-based KD that minimize forward Kullback-Leibler divergence (KLD) in free-run generation. They find that such approaches can lead the student model to overestimate the low-probability regions of the teacher’s distribution. To address this issue, MINILLM proposes replacing the forward KLD objective (as shown in Eq. 7) with reverse KLD, i.e., $reverse~{}\mathrm{KLD}:=KL(p^{\bf s}||p^{\bf t})$, which is better suited for knowledge distillation on generative large language models (LLMs). GKD [99] identifies a challenge in distilling knowledge into auto-regressive student models, which is the train-test distribution mismatch. Specifically, partial sequences encountered by the student during generation phase can be significantly different from the ones observed during the training phase. To address this issue, GKD trains the student on its self-generated output sequences by leveraging feedback (logits) from the teacher on such sequences. It also facilitates the integration of distillation with RL fine-tuning of LLMs. Huang et al. [100] further combines the hint-based KD with multitask in-context learning.
@@
Hint-based KD. Sun et al. [101] proposes patient knowledge distillation to help shallow BERT student learn from multiple intermediate features [102, 103, 104, 25] of the deep BERT teacher. Li et al. [105] utilizes hints extracted from both intermediate hidden states and attention distributions to enhance the training of Non-AutoRegressive Translation (NART) models. The distilled NART models attain performance similar to a powerful LSTM-based AutoRegressive Translation (ART) baseline in various machine translation tasks. Notably, the distilled NART models exhibit a speed improvement of 17$\times$ over their ART counterparts. Similarly, Mukherjee et al. [106] employs MBERT [107] as a teacher model to guide the training of smaller BiLSTM models, resulting in a significant 51$\times$ speed improvement for batch inference. MobileBERT [103] implements two objective functions to distill knowledge from a BERT teacher incorporated with inverted bottleneck, including attention distributions and hidden states, to a slimmed-down version of BERT as the student model. In addition to these objectives, TinyBERT [104] extends the distillation process by also transferring knowledge from the teacher’s logits to the student model. Wang et al. [25] proposes MINILM to train the student networks to mimic both the attention distributions and the scaled dot-product between values of teacher models. It retains more than 99% accuracy on General Language Understanding Evaluation (GLUE) benchmark tasks using 50% of the Transformer parameters and computations of the teacher model.
@@
Several recent studies [110, 111, 112, 113, 114, 7, 115, 8, 116] have showed the promising results in fine-tuning small models on the outputs generated by LLMs. Specifically, they first prompt a very large teacher model, such as GPT-3 with 175B parameters, to solve complex questions via chain-of-thought (CoT) reasoning [117]. Then, they use the generated instructions [7, 115], explanations [8] or rationales [110] to fine-tune a much smaller student. Furthermore, Zhu et al. [118] proposes PaD, which utilizes program-aided reasoning, such as an additional Python interpreter, to help student models overcome faulty steps in the CoT reasoning with automated error checking. Shridhar et al. [119] introduces Socratic CoT that trains a combination of two small distilled models: a problem decomposer and a subproblem solver. These models work together to decompose and solve complex problems effectively. Fu et al. [120] demonstrates that specializing the small student model’s abilities for the specific reasoning task, by transferring the knowledge in larger teacher model’s generic directions, yields promising results. Jiang et al. [7] identifies a shortcoming in the aforementioned approaches, as they lack the ability to incorporate feedback for identifying challenging instructions where the student model’s performance falls short. To address this limitation, they propose a three-stage adversarial loop of distillation that incorporates feedback and addresses these challenging instructions.
Other KDs. Xu et al. [121] compresses the BERT model by progressive module replacing. The original model is first divided into multiple modules, and more compact substitutes are created for these modules. Then, they randomly substitute the original modules with their corresponding substitutes and train the compact modules without any additional loss functions.
Jha et al. [122] discovers that fitting both the student and teacher models, each with numerous parameters, into GPU memory with traditional knowledge distillation is impractical. To overcome this limitation, they propose a teacher-free task-agnostic distillation method which uses a truncated version (student) of the large model (teacher) for initialization [123] and then continuing the pre-training of student, without employing any distillation loss.
@@ 3.2.3 KD for Transformer-Based Vision Models
Logits-based KD. While conventional KD in vision tasks typically involves using soft labels [86, 125, 126] generated by a powerful teacher network, Touvron et al. [127] makes a noteworthy discovery. Their experiments on the ImageNet benchmark reveal that employing hard labels, which are essentially the maximum of the score values predicted by the teacher model, leads to superior results. Additionally, they introduce a novel distillation token into vision Transformers (ViTs)[128]. TinyViT [27] observes that smaller ViTs can benefit from larger teacher using massive pre-training data, for example, pre-training (distilling) student on ImageNet-21k while fine-tuning student on ImageNet-1k. And they propose TinyViT to save computational memory by storing data augmentation information and logits in advance for large teacher models. Ren et al. [129] argues that student accuracy is primarily influenced by the inductive bias of the teacher models rather than their accuracy. To address this, they introduce a cross-inductive bias distillation approach, which involves distilling the student model with multiple teachers exhibiting distinct architectural inductive biases. To enhance the utilization of the inductive bias in CNNs and offer more image-friendly guidance, CSKD [130] removes the pooling operation after the final feature map and treats the features at each position as individual samples. Subsequently, these position-based features are fed into the classifier, leading to the generation of dense predictions for each corresponding position. The hard labels derived from the CNN’s dense predictions are used as target labels, and cross-entropy serves as the loss function for spatial knowledge transfer.
Hint-based KD. Chen et al. [131] introduces a two-stage DearKD framework. In the initial stage, knowledge distillation is employed from the intermediate layers of the CNN to help Transformer-based student capture the inductive biases. Subsequently, the second stage trains the student without distillation. Hao et al. [28] proposes a fine-grained manifold distillation method to fully utilize the patch-level and batch-level information inside Transformer-based architectures. Specifically, they manually select intermediate layers from teacher-student pair and distilled the student via decoupled relation maps. ViTKD et al. [132] delves into the characteristics of feature maps in ViTs and formulated three practical guidelines for hint-based distillation in ViTs, namely: (i) generating hints is more effective than mimicking in deeper layers, (ii) distillation on shallower layers remains suitable for ViTs distilled through mimicking, and (iii) features from the feed-forward network (FFN) are better than those from the multi-head attention (MHA) for the distillation process.
@@ 3.2.4 Discussion
One of the strengths of ViT models is their ability to scale to high parametric complexity, but this requires substantial computational resources and comes at a high cost. KD can be used to transfer knowledge into more compact student models, but there are still challenges in the context of vision that require further research. The first challenge is related to training costs. Both logits-based and hint-based KD methods require large GPU memory for the distillation process. Vision tasks differ from NLP as the input for vision is in the form of images, which are significantly larger in size compared to the limited sequence length used in NLP tasks. During distillation, even though the teacher network doesn’t require backward propagation, each forward pass consumes a significant amount of GPU memory due to the activations of intermediate features. Therefore, finding ways to reduce training costs and shorten training times is an area that needs exploration. Additionally, choosing the most suitable teacher model for Transformer-based students is an open question. For instance, in the context of ViTs, determining whether a CNN or a Transformer teacher is more effective for the student model is a potential research avenue. Moreover, as proposed by Hao et al. [126], there is a necessity for designing and evaluating KD approaches within practical scenarios, moving away from the limitations of small-scale datasets. This ensures that KD methods are applicable and effective in real-world, large-scale settings.